# On the Parameter Identifiability of Partially Observed Linear Causal Models

### v1

@@ Abstract
Linear causal models are important tools for modeling causal dependencies and yet in practice, only a subset of the variables can be observed. In this paper, we examine the parameter identifiability of these models by investigating whether the edge coefficients can be recovered given the causal structure and partially observed data. Our setting is more general than that of prior research—we allow all variables, including both observed and latent ones, to be flexibly related, and we consider the coefficients of all edges, whereas most existing works focus only on the edges between observed variables. Theoretically, we identify three types of indeterminacy for the parameters in partially observed linear causal models. We then provide graphical conditions that are sufficient for all parameters to be identifiable and show that some of them are provably necessary. Methodologically, we propose a novel likelihood-based parameter estimation method that addresses the variance indeterminacy of latent variables in a specific way and can asymptotically recover the underlying parameters up to trivial indeterminacy. Empirical studies on both synthetic and real-world datasets validate our identifiability theory and the effectiveness of the proposed method in the finite-sample regime.
@@ 1 Introduction and Related Work
In this paper, we focus on the the identification of parameters of a partially observed model. Specifically, given the causal structure of and observational data from a partially observed causal model, we are interested in identifying all the parameters, and thus the underlying causal model can be fully specified. To identify the parameters, a classical way is to project the directed acyclic graph (DAG) with latent variables to an acyclic directed mixed graph (ADMG) or partially ancestral graph [40], without explicitly modeling the latent confounders. Based on ADMG, graphical criteria such as half-trek [19], G-criterion [9], and some further developments [50, 28] have been proposed to establish the parameter identifiability. Another way is to leverage do-calculus, proxy variables, and instrumental variables [45, 37, 24] to identify the direct causal effect, which corresponds to the edge coefficient in linear causal models. For a more detailed discussion of related work, please refer to Appendix C.
@@ 2.1 Problem Setting
Let $\mathcal{G}:=(\mathbf{V}_{\mathcal{G}},\mathbf{E}_{\mathcal{G}})$ be a DAG. Each variable $V_{i}\in\mathbf{V}_{\mathcal{G}}$ follows a linear structural equation model $\mathsf{V}_{i}=\sum\nolimits_{\mathsf{V}_{j}\in\text{Pa}_{\mathcal{G}}(\mathsf{V}_{i})}f_{ji}\mathsf{V}_{j}+\epsilon_{\mathsf{V}_{i}}$, where $\mathbf{V}_{\mathcal{G}}:=\mathbf{L}_{\mathcal{G}}\cup\mathbf{X}_{\mathcal{G}}=\{\mathsf{L}_{i}\}_{i=1}^{m}\cup\{\mathsf{X}_{i}\}_{i=m+1}^{m+n}$ contains $m$ latent variables and $n$ observed variables. $\text{Pa}_{\mathcal{G}}(\mathsf{V}_{i})$ denotes the parent set of $\mathsf{V}_{i}$, $f_{ji}$ denotes the edge coefficient from $V_{j}$ to $V_{i}$, and $\epsilon_{\mathsf{V}_{i}}$ represents the Gaussian noise term of $\mathsf{V}_{i}$. We drop the subscript $\mathcal{G}$ in $\mathbf{L}_{\mathcal{G}}$ and $\mathbf{X}_{\mathcal{G}}$ when the context is clear.
We use $\mathsf{V}$, $\mathbf{V}$, and $\mathcal{V}$ to denote a random variable, a set of variables, and a set of sets of variables, respectively. In Definition 1, the relations between variables can also be written in the matrix form as $\mathbf{V}_{\mathcal{G}}=F^{T}\mathbf{V}_{\mathcal{G}}+\epsilon_{\mathbf{V}_{\mathcal{G}}}$, where $F=(f_{ji})_{i,j\in[d]}$ is the weighted adjacency matrix. Here, $f_{ji}\neq 0$ if and only if $V_{j}$ is a parent of $V_{i}$ in $\mathcal{G}$. We also write
@@
where $\Omega$ is the diagonal covariance matrix of $\epsilon_{\mathbf{V}_{\mathcal{G}}}$. Our objective is to identify $F$, the causal edge coefficient of the model, given observational data and the causal structure $\mathcal{G}$. Note that variables are partially observed and thus we only have access to i.i.d. samples of observed variables $\mathbf{X}_{\mathcal{G}}$. As variables are jointly Gaussian, asymptotically our observation can be summarized as population covariance over observed variables, i.e., ${\Sigma}_{\mathbf{X}_{\mathcal{G}}}$. In other words, we aim to identify $F$, given ${\Sigma}_{\mathbf{X}_{\mathcal{G}}}$ and $\mathcal{G}$. The identification of parameters is important in that, once we identify the parameters, the underlying causal model is fully specified, and thus we can flexibly calculate causal effects, infer interventional distributions, and finally answer counterfactual queries [37]. It is worth noting that, for parameter identification, the structure $\mathcal{G}$ is assumed to be known, which is different from the setting of causal discovery where the goal is to identify $\mathcal{G}$ from data.
@@
Figure 1: Examples to illustrate the advantage of our framework. Within our framework, it can be shown that $\mathcal{G}_{1}$’s parameters can be fully identified while $\mathcal{G}_{2}$’s cannot. In contrast, the latent projection framework cannot even differentiate $\mathcal{G}_{1}$ from $\mathcal{G}_{2}$ as they share the same ADMG (c) after projection. Furthermore, with ADMG, any edge coefficient that involves a latent variable cannot be considered.
@@ 2.2 Framework Comparison
When there is no latent variable, it has been shown that all the parameters are identifiable [8]. However, the problem becomes very challenging when latent variables exist. There are two lines of research for this problem. One line of thought focuses on the use of do-calculus, proxy variables, and instrumental variables to identify direct causal effects among observed variables [45, 37, 24] (in linear models the direct causal effect is captured by the edge coefficient). Another line of thought addresses latent confounders by projecting a DAG with latent variables into an ADMG, where the confounding effects of latent variables are simplified and represented by correlation among noise terms [19, 9, 50, 28]. An example is in Figure 1, where (a) is the original graph and (c) is the projected ADMG whose bidirected edges correspond to correlated noise terms.
@@ 3.1 Definition of Parameter Identifiability and Indeterminacy
Let $\theta=(F,\Omega)\in\Theta$. We say that $F$ is generically identifiable, if the mapping $\phi(\theta)=\Sigma_{\mathbf{X}_{\mathcal{G}}}$ is injective, for almost all $\theta\in\Theta$ with respect to the Lebesgue measure.
@@
Definition 2 indicates if $\theta$ is identifiable, then there is not $\theta^{\prime}\in\Theta$ such that $\theta^{\prime}$ entails the same observation as that of $\theta$. As in the typical literature of parameter identification, we consider generic identifiability to rule out some rare cases where the parameters for that structure is generally identifiable, but with some specific parameterization, the parameters cannot be identified. This is similar to faithfulness in causal discovery [48] and we will provide an example in Example 1. We next introduce three important indeterminacies about parameter identification when latent variable exists.
Lemma 1 (Indeterminacy of Variance of Latent Variables).
We can rescale any latent variables and change the coefficients in specific ways such that the entailed observation $\Sigma_{\mathbf{X}_{\mathcal{G}}}$ is unchanged. In other words, the variance of latent variables can never be identified, or equivalently, the variance of the noise terms of latent variables never be identified.
The crucial implication of Lemma 1 is that, we have to assume all latent variables have a unit variance; otherwise parameters involving latent variables can never be identified, due to the indeterminacy of the variance of latent variables. The proof of Lemma 1 can be found in Appendix. We note that a similar lemma is provided in [4], and yet the setting we considered is more general and takes that of [4] as a special case. Concretely, in our setting, latent variables and observed variables can be related in a flexible fashion – latent variables can be the cause or the effect of observed variables, while in [4] they only consider the scenario where latent variables cannot be the effect of observed variables.
@@
Definition 3 (Group Sign Indeterminacy).
If there exists a sign matrix $S\in\{-1,1\}^{(m+n)\times(m+n)}$ with at least one entry of $-1$ such that $(S\odot F,\Omega)$ entails the same observation as that of $(F,\Omega)$, then we say that there exists a group sign indeterminacy in the identification of parameters.
@@
In Figure 2 (b), given the structure and $\Sigma_{\mathbf{X}_{\mathcal{G}}}$, by assuming unit variance, the parameters are generally identifiable up to group sign indeterminacy. Specifically, there exist three equality constraints with three free parameters: $f_{1,2}f_{1,3}=\sigma_{2,3}$, $f_{1,2}f_{1,4}=\sigma_{2,4}$, and $f_{1,3}f_{1,4}=\sigma_{3,4}$. The solutions are: (i) $f_{1,2}=\sqrt{\frac{\sigma_{2,3}\sigma_{2,4}}{\sigma_{3,4}}}$, $f_{1,3}=\sigma_{2,3}/f_{1,2}$, $f_{1,4}=\sigma_{2,4}/f_{1,2}$ and (ii) $f_{1,2}=-\sqrt{\frac{\sigma_{2,3}\sigma_{2,4}}{\sigma_{3,4}}}$, $f_{1,3}=\sigma_{2,3}/f_{1,2}$, $f_{1,4}=\sigma_{2,4}/f_{1,2}$. The two solutions are different only in terms of group sign. However, if we set $f_{1,2}=0$, then the parameters are not identifiable. These rare cases of parameters are of zero Lebesgue measure so we rule out these cases for the definition of identifiability, as in Definition 2.
We note that, as the indeterminacy of group sign is rather minor, in the following if the parameters are identifiable only upto group sign indeterminacy, we still say that the parameters are identifiable.
Definition 4 (Orthogonal Transformation Indeterminacy, following factor analysis [43, 7]).
If there exists a non-diagonal orthogonal matrix $Q$ such that $(QF,\Omega)$ entails the same observation as that of $(F,\Omega)$ and $QF$ and $F$ share the same support, then there exists an orthogonal transformation indeterminacy in the identification of parameters.
The orthogonal transformation indeterminacy is the major indeterminacy we consider in the presence of latent variables. Here we only give the definition and will later provide Theorem 2 with an example that captures the scenarios where such indetermincay exists.

### v2

@@ Abstract
Linear causal models are important tools for modeling causal dependencies and yet in practice, only a subset of the variables can be observed. In this paper, we examine the parameter identifiability of these models by investigating whether the edge coefficients can be recovered given the causal structure and partially observed data. Our setting is more general than that of prior research—we allow all variables, including both observed and latent ones, to be flexibly related, and we consider the coefficients of all edges, whereas most existing works focus only on the edges between observed variables. Theoretically, we identify three types of indeterminacy for the parameters in partially observed linear causal models. We then provide graphical conditions that are sufficient for all parameters to be identifiable and show that some of them are provably necessary. Methodologically, we propose a novel likelihood-based parameter estimation method that addresses the variance indeterminacy in a specific way and can asymptotically recover the underlying parameters up to trivial indeterminacy. Empirical studies on both synthetic and real-world datasets validate our identifiability theory and the effectiveness of the proposed method in the finite-sample regime. Code: https://github.com/dongxinshuai/scm-identify.
@@ 1 Introduction and Related Work
In this paper, we focus on the the identification of parameters of a partially observed model. Specifically, given the causal structure of and observational data from a partially observed causal model, we are interested in identifying all the parameters, and thus the underlying causal model can be fully specified. To identify the parameters, a classical way is to project the directed acyclic graph (DAG) with latent variables to an acyclic directed mixed graph (ADMG) or partially ancestral graph [42], without explicitly modeling the latent confounders. Based on ADMG, graphical criteria such as half-trek [20], G-criterion [9], and some further developments [51, 29] have been proposed to establish the parameter identifiability. Another way is to leverage do-calculus, proxy variables, and instrumental variables [47, 39, 25] to identify the direct causal effect, which corresponds to the edge coefficient in linear causal models. For a more detailed discussion of related work, please refer to Appendix D.
@@ 2.1 Problem Setting
Let $\mathcal{G}:=(\mathbf{V}_{\mathcal{G}},\mathbf{E}_{\mathcal{G}})$ be a DAG. Each variable $V_{i}\in\mathbf{V}_{\mathcal{G}}$ follows a linear structural equation model $\mathsf{V}_{i}=\sum\nolimits_{\mathsf{V}_{j}\in\text{Pa}_{\mathcal{G}}(\mathsf{V}_{i})}f_{j,i}\mathsf{V}_{j}+\epsilon_{\mathsf{V}_{i}}$, where $\mathbf{V}_{\mathcal{G}}:=\mathbf{L}_{\mathcal{G}}\cup\mathbf{X}_{\mathcal{G}}=\{\mathsf{L}_{i}\}_{i=1}^{m}\cup\{\mathsf{X}_{i}\}_{i=m+1}^{m+n}$ contains $m$ latent variables and $n$ observed variables. $\text{Pa}_{\mathcal{G}}(\mathsf{V}_{i})$ denotes the parent set of $\mathsf{V}_{i}$, $f_{j,i}$ denotes the edge coefficient from $V_{j}$ to $V_{i}$, and $\epsilon_{\mathsf{V}_{i}}$ represents the Gaussian noise term of $\mathsf{V}_{i}$.
We drop the subscript $\mathcal{G}$ in $\mathbf{L}_{\mathcal{G}}$ and $\mathbf{X}_{\mathcal{G}}$ when the context is clear. We use $\mathsf{V}$, $\mathbf{V}$, and $\mathcal{V}$ to denote a random variable, a set of variables, and a set of sets of variables, respectively. In Definition 1, the relations between variables can also be written in the matrix form as $\mathbf{V}_{\mathcal{G}}=F^{T}\mathbf{V}_{\mathcal{G}}+\epsilon_{\mathbf{V}_{\mathcal{G}}}$, where $F=(f_{j,i})_{i,j\in[m+n]}$ is the weighted adjacency matrix. Here, $f_{j,i}\neq 0$ if and only if $V_{j}$ is a parent of $V_{i}$ in $\mathcal{G}$. We also write
@@
where $\Omega$ is the diagonal covariance matrix of $\epsilon_{\mathbf{V}_{\mathcal{G}}}$.
Our objective is to identify $F$, the causal edge coefficients of the model, given observational data and the causal structure $\mathcal{G}$. Denote by $\Sigma_{\mathbf{L}}$ and $\Sigma_{\mathbf{X}}$ the population covariance matrix of latent variables $\mathbf{L}$ and observed variables $\mathbf{X}$, respectively; their precise formulations are provided in Proposition 1. We also denote by $\sigma_{i,j}$ the $(i,j)$-th entry of $\Sigma_{\mathbf{X}}$. In this work, we assume that the noise terms of latent variables, $\epsilon_{\mathbf{L}}$, have unit variance, i.e., $\Omega_{\epsilon_{\mathbf{L}}}=I$, which will be justified later in Section 3.1. Note that variables are partially observed and thus we only have access to i.i.d. samples of observed variables $\mathbf{X}$. As variables are jointly Gaussian, the observations can asymptotically be summarized as the population covariance matrix ${\Sigma}_{\mathbf{X}}$. In other words, we aim to identify $F$ and $\Omega$ given ${\Sigma}_{\mathbf{X}}$ and $\mathcal{G}$. The identification of parameters is important in that, once we identify the parameters, the underlying causal model is fully specified, and thus we can flexibly calculate causal effects, infer interventional distributions, and finally answer counterfactual queries [39]. It is worth noting that, for parameter identification, the structure $\mathcal{G}$ is assumed to be known, which is different from the setting of causal discovery where the goal is to identify $\mathcal{G}$ from data.
@@
Figure 1: Illustrations of the advantage of our framework. Within our framework, it can be shown that $\mathcal{G}_{1}$’s parameters can be identified (up to sign) while $\mathcal{G}_{2}$’s cannot. In contrast, the latent projection framework cannot even differentiate $\mathcal{G}_{1}$ from $\mathcal{G}_{2}$ as they share the same ADMG (c) after projection. Furthermore, with ADMG, any edge coefficient that involves a latent variable cannot be considered.
@@ 2.2 Framework Comparison
Without latent variables, it has been shown all parameters are identifiable [8]. However, the problem becomes very challenging when latent variables exist. There are two lines of research. One focuses on the use of do-calculus, proxy variables, and instrumental variables to identify direct causal effects among observed variables [47, 39, 25] (in linear models the direct causal effect is captured by the edge coefficient). Another line addresses latent confounders by projecting a DAG with latent variables into an ADMG, where the confounding effects of latent variables are simplified and represented by correlation among noise terms [20, 9, 51, 29]. An example is in Figure 1, where (a) is the original graph and (c) is the projected ADMG whose bidirected edges correspond to correlated noise terms.
@@ 3.1 Definition of Parameter Identifiability and Indeterminacy
Let $\theta=(F,\Omega)\in\Theta$. We say that $\theta$ is generically identifiable, if the mapping $\phi(\theta)=\Sigma_{\mathbf{X}}$ is injective, for almost all $\theta\in\Theta$ with respect to the Lebesgue measure.
@@
Definition 2 indicates if parameter $\theta$ is identifiable, then there does not exist $\theta^{\prime}\in\Theta$ that entails the same observations as those of $\theta$. As in the typical literature of parameter identification, we consider generic identifiability to rule out some rare cases where the parameters for that structure is generally identifiable, but with some specific parameterization, the parameters cannot be identified. This is similar to faithfulness in causal discovery [49] and we will provide an example in Example 1. We next introduce three important indeterminacies about parameter identification when latent variables exist.
Theorem 1 (Indeterminacy of Scaling of $\Omega_{\epsilon_{\mathbf{L}}}$).
Consider a model that follows Def. 1 with number of latent variables $m\geq 1$ and $\theta=(F_{\mathbf{L}\mathbf{L}},F_{\mathbf{L}\mathbf{X}},F_{\mathbf{X}\mathbf{L}},F_{\mathbf{X}\mathbf{X}},\Omega_{\epsilon_{\mathbf{L}}},\Omega_{\epsilon_{\mathbf{X}}})$. Let $\Lambda$ be any invertible diagonal matrix, and $\tilde{\theta}=(\tilde{F}_{\mathbf{L}\mathbf{L}},\tilde{F}_{\mathbf{L}\mathbf{X}},\tilde{F}_{\mathbf{X}\mathbf{L}},\tilde{F}_{\mathbf{X}\mathbf{X}},\tilde{\Omega}_{\epsilon_{\mathbf{L}}},\tilde{\Omega}_{\epsilon_{\mathbf{X}}})$, where
$\tilde{F}_{\mathbf{L}\mathbf{L}}=\Lambda^{-1}{F}_{\mathbf{L}\mathbf{L}}\Lambda,~{}\tilde{F}_{\mathbf{L}\mathbf{X}}=\Lambda^{-1}{F}_{\mathbf{L}\mathbf{X}},~{}\tilde{F}_{\mathbf{X}\mathbf{L}}={F}_{\mathbf{X}\mathbf{L}}\Lambda,~{}\tilde{F}_{\mathbf{X}\mathbf{X}}={F}_{\mathbf{X}\mathbf{X}},~{}\tilde{\Omega}_{\epsilon_{\mathbf{L}}}=\Lambda^{2}\Omega_{\epsilon_{\mathbf{L}}},~{}\tilde{\Omega}_{\epsilon_{\mathbf{X}}}=\Omega_{\epsilon_{\mathbf{X}}}.$
Then, $\tilde{\theta}$ and $\theta$ entail the same observations, i.e., $\tilde{\Sigma}_{\mathbf{X}}=\Sigma_{\mathbf{X}}$. Furthermore, we have $\tilde{\Sigma}_{\mathbf{L}}=\Lambda\Sigma_{\mathbf{L}}\Lambda$.
A similar theoretical result is provided in [4], and yet our setting is much more general and takes that of [4] as a special case: in our setting, all variables including latent and observed ones can be arbitrarily related while in [4] latent variables cannot be the effect of observed variables.
Remark 1 (Implication of Theorem 1).
A key implication of Theorem 1 is that, without further assumption, the edge coefficients involving latent variables, i.e., $({F}_{\mathbf{L}\mathbf{L}},{F}_{\mathbf{L}\mathbf{X}},{F}_{\mathbf{X}\mathbf{L}})$, can never be identified, as there always exists a diagonal matrix $\Lambda$ such that $\tilde{\theta}$ and $\theta$ entail the same observations but $(\tilde{F}_{\mathbf{L}\mathbf{L}},\tilde{F}_{\mathbf{L}\mathbf{X}},\tilde{F}_{\mathbf{X}\mathbf{L}})\neq({F}_{\mathbf{L}\mathbf{L}},{F}_{\mathbf{L}\mathbf{X}},{F}_{\mathbf{X}\mathbf{L}})$. Thus, in the rest of this paper, we assume that the noise terms of latent variables, $\epsilon_{\mathbf{L}}$, have unit variance, i.e., $\Omega_{\epsilon_{\mathbf{L}}}=I$. Under this assumption, we have $(\tilde{\Omega}_{\epsilon_{\mathbf{L}}})_{i,i}=\Lambda_{i,i}^{2}(\Omega_{\epsilon_{\mathbf{L}}})_{i,i}=1,i\in[m]$, which implies $\Lambda_{i,i}=\pm 1$. As such, this assumption makes parameter identifiability possible. However, even though we fix the scaling of $\Omega_{\epsilon_{\mathbf{L}}}$, there still exists indeterminacy about the sign of parameters, captured by Theorem 2.
Theorem 2 (Group Sign Indeterminacy).
Consider a model that follows Def. 1 with number of latent variables $m\geq 1$, $\theta=(F_{\mathbf{L}\mathbf{L}},F_{\mathbf{L}\mathbf{X}},F_{\mathbf{X}\mathbf{L}},F_{\mathbf{X}\mathbf{X}},\Omega_{\epsilon_{\mathbf{L}}},\Omega_{\epsilon_{\mathbf{X}}})$, and $\Omega_{\epsilon_{\mathbf{L}}}=I$. Let $S$ be a diagonal sign matrix (entries are either $1$ or $-1$), and $\tilde{\theta}=(\tilde{F}_{\mathbf{L}\mathbf{L}},\tilde{F}_{\mathbf{L}\mathbf{X}},\tilde{F}_{\mathbf{X}\mathbf{L}},\tilde{F}_{\mathbf{X}\mathbf{X}},\tilde{\Omega}_{\epsilon_{\mathbf{L}}},\tilde{\Omega}_{\epsilon_{\mathbf{X}}})$, where
$\tilde{F}_{\mathbf{L}\mathbf{L}}=S{F}_{\mathbf{L}\mathbf{L}}S,~{}\tilde{F}_{\mathbf{L}\mathbf{X}}=S{F}_{\mathbf{L}\mathbf{X}},~{}\tilde{F}_{\mathbf{X}\mathbf{L}}={F}_{\mathbf{X}\mathbf{L}}S,~{}\tilde{F}_{\mathbf{X}\mathbf{X}}={F}_{\mathbf{X}\mathbf{X}},~{}\tilde{\Omega}_{\epsilon_{\mathbf{L}}}=\Omega_{\epsilon_{\mathbf{L}}}=I,~{}\tilde{\Omega}_{\epsilon_{\mathbf{X}}}=\Omega_{\epsilon_{\mathbf{X}}}.$
Then, $\tilde{\theta}$ and $\theta$ entail the same observations, i.e., $\tilde{\Sigma}_{\mathbf{X}}=\Sigma_{\mathbf{X}}$, and $(\tilde{\Sigma}_{\mathbf{L}})_{ii}=(\Sigma_{\mathbf{L}})_{ii},~{}\forall i\in[m]$.
@@
Remark 2 (Remark on Theorem 2).
The indeterminacy described in Theorem 2 is referred to as group sign indeterminacy for the following reason: According to the theorem, flipping the sign of $S_{i,i}$ is equivalent to flipping the signs of all coefficients of edges involving the latent variable $\mathsf{L}_{i}$. This transformation preserves the resulting observations $\Sigma_{\mathbf{X}}$. In essence, each group consists of coefficients of edges involving a particular latent variable.
@@
In Figure 2 (b), given the structure and $\Sigma_{\mathbf{X}}$, by assuming $\Omega_{\epsilon_{\mathbf{L}}}=I$, the parameters are generally identifiable up to group sign indeterminacy. Specifically, there exist three equality constraints with three free parameters: $f_{1,2}f_{1,3}=\sigma_{2,3}$, $f_{1,2}f_{1,4}=\sigma_{2,4}$, and $f_{1,3}f_{1,4}=\sigma_{3,4}$. The solutions are: (i) $f_{1,2}=\sqrt{\frac{\sigma_{2,3}\sigma_{2,4}}{\sigma_{3,4}}}$, $f_{1,3}=\sigma_{2,3}/f_{1,2}$, $f_{1,4}=\sigma_{2,4}/f_{1,2}$ and (ii) $f_{1,2}=-\sqrt{\frac{\sigma_{2,3}\sigma_{2,4}}{\sigma_{3,4}}}$, $f_{1,3}=-\sigma_{2,3}/f_{1,2}$, $f_{1,4}=-\sigma_{2,4}/f_{1,2}$. The two solutions are different only in terms of group sign. However, if we set $f_{1,2}=0$, then the parameters are not identifiable (as we will encounter division where the divisor is zero). These rare cases of parameters are of zero Lebesgue measure so we rule out these cases for the definition of identifiability, as in Definition 2.
Intuitively speaking, group sign indeterminacy arises because one may multiply the latent variable $\mathsf{L}_{i}$ by $-1$ and accordingly flip the signs of all edge coefficients involving $\mathsf{L}_{i}$. Note that such an indeterminacy is rather minor for the following reason. (i) In practice, we can always anchor the sign of some edges according to our preference or prior knowledge in order to eliminate the group sign indeterminacy. For example, in Figure 4, if we expect that L2 should be understood as Extraversion instead of non-Exterversion, we can add one additional constraint during our parameter estimation such that the edge coefficient from L2 to E1 ("I am the life of party.") will be positive (as we believe E1 should be positively related to Extraversion). (ii) On the other hand, there are some application scenarios that are not influenced by the group sign indeterminacy, such as causal effect estimations between certain variables. We note that, as the indeterminacy of group sign is rather minor, in the following if the parameters are identifiable only up to group sign indeterminacy, we still say that the parameters are identifiable.
Definition 3 (Orthogonal Transformation Indeterminacy).
Consider a model that follows Def. 1 with number of latent variables $m\geq 1$, $\theta=(F_{\mathbf{L}\mathbf{L}},F_{\mathbf{L}\mathbf{X}},F_{\mathbf{X}\mathbf{L}},F_{\mathbf{X}\mathbf{X}},\Omega_{\epsilon_{\mathbf{L}}},\Omega_{\epsilon_{\mathbf{X}}})$, and $\Omega_{\epsilon_{\mathbf{L}}}=I$. We say that there exists an orthogonal transformation indeterminacy in the identification of parameters if there exists a non-diagonal orthogonal matrix $Q$ such that $(F_{\mathbf{L}\mathbf{L}},F_{\mathbf{L}\mathbf{X}},F_{\mathbf{X}\mathbf{L}},F_{\mathbf{X}\mathbf{X}},\Omega_{\epsilon_{\mathbf{L}}},\Omega_{\epsilon_{\mathbf{X}}})$ and $(\tilde{F}_{\mathbf{L}\mathbf{L}},\tilde{F}_{\mathbf{L}\mathbf{X}},\tilde{F}_{\mathbf{X}\mathbf{L}},\tilde{F}_{\mathbf{X}\mathbf{X}},\tilde{\Omega}_{\epsilon_{\mathbf{L}}},\tilde{\Omega}_{\epsilon_{\mathbf{X}}})$ share the same support and entail the same observations, where
$\tilde{F}_{\mathbf{L}\mathbf{L}}=Q^{T}{F}_{\mathbf{L}\mathbf{L}}Q,~{}\tilde{F}_{\mathbf{L}\mathbf{X}}=Q^{T}{F}_{\mathbf{L}\mathbf{X}},~{}\tilde{F}_{\mathbf{X}\mathbf{L}}={F}_{\mathbf{X}\mathbf{L}}Q,~{}\tilde{F}_{\mathbf{X}\mathbf{X}}={F}_{\mathbf{X}\mathbf{X}},~{}\tilde{\Omega}_{\epsilon_{\mathbf{L}}}=\Omega_{\epsilon_{\mathbf{L}}}=I,~{}\tilde{\Omega}_{\epsilon_{\mathbf{X}}}=\Omega_{\epsilon_{\mathbf{X}}}.$
The orthogonal transformation indeterminacy is the major indeterminacy we consider in the presence of latent variables. Such an indeterminacy also arises in factor analysis [45, 7], which can be viewed as a special case of the data generating procedure considered in Definition 1. Here we only give the definition and will later provide Theorem 4 with an example that captures the scenarios where such indeterminacy exists.